# A Framework for Supervised and Unsupervised Segmentation and Classification of Materials Microstructure Images

### v1

@@ Abstract
Microstructure of materials is often characterized through image analysis to understand processing-structure-properties linkages. We propose a largely automated framework that integrates unsupervised and supervised learning methods to classify micrographs according to microstructure phase/class and, for multiphase microstructures, segments them into different homogeneous regions. With the advance of manufacturing and imaging techniques, the ultra-high resolution of imaging that reveals the complexity of microstructures and the rapidly increasing quantity of images (i.e., micrographs) enables and necessitates a more powerful and automated framework to extract materials characteristics and knowledge. The framework we propose can be used to gradually build a database of microstructure classes relevant to a particular process or group of materials, which can help in analyzing and discovering/identifying new materials. The framework has three steps: (1) segmentation of multiphase micrographs through a recently developed score-based method so that different microstructure homogeneous regions can be identified in an unsupervised manner; (2) identification and classification of homogeneous regions of micrographs through an uncertainty-aware supervised classification network trained using the segmented micrographs from Step $1$ with their identified labels verified via the built-in uncertainty quantification and minimal human inspection; (3) supervised segmentation (more powerful than the segmentation in Step $1$) of multiphase microstructures through a segmentation network trained with micrographs and the results from Steps $1$-$2$ using a form of data augmentation. This framework can iteratively characterize/segment new homogeneous or multiphase materials while expanding the database to enhance performance. The framework is demonstrated on various sets of materials and texture images.
@@ 1 Motivation and Introduction
Microstructure plays a vital role in determining physical properties of materials and understanding the relations among the processes, structures, and properties DeCost and Holm, (2015). Recently, analyzing images of multiphase materials, often called micrographs O’Mara et al., (2016); Puchala et al., (2016); Michel and Meredig, (2016), has attracted much attention across different materials fields (e.g., metals Lewandowski and Seifi, (2016), polymer composites Ligon et al., (2017), and ceramics Chen et al., (2019b); Moritz and Maleksaeedi, (2018), etc.) due to burgeoning manufacturing and imaging techniques Balla et al., (2012); Bandyopadhyay and Heer, (2018). For example, the ultra-high carbon steel (UHCS) data set DeCost et al., (2017b) includes $961$ SEM micrographs, of which two multiphase microstructures are shown in Fig. 1(a) and 1(b). Fig. 1(a) is a micrograph containing a single spheroidite phase, while Fig. 1(b) contains ferritic, proeutectoid, spheroidite, and Widmanstätten phases. Because the phase characteristics impact material properties (e.g., toughness and crack initiation etc.) Hecht et al., (2016); Carter et al., (2014), identifying and classifying the individual phases present in a micrograph is an important step towards understanding the physical properties of the material. Throughout this paper, by “multiphase” we mean a microstructure whose stochastic nature (different phases have different stochastic nature) varies spatially, sometimes referred to as microstructure nonstationarity. This includes certain steels that have different phases present simultaneously by design, as well as certain composites whose stochastic microstructure composition varies spatially due to variation in the processing conditions. By “single-phase”, we mean a homogeneous microstructure whose stochastic nature is spatially stationary. Throughout the paper, the term “microstructure class” (MC) refers to the label assigned to each homogeneous region (HR) of connected pixels within a micrograph with each HR comprised of a single phase.
@@
Figure 1: Examples of single phase (a) and multiphase materials (b)-(f)00footnotemark: 0. (a)-(b) Two SEM images of the UHCS data set DeCost et al., (2017b); Hecht et al., (2016). Notice that the two images are not in the same magnification, a commonly encountered issue of data quality in real data sets. (c) A X-ray CT (XCT) image of dendrite growth from a solidification experiment on a Al-Zn alloy Stan et al., (2020). (d) A serial sectioning (SS) image of dendrite growth from a coarsening experiment on a Pb-Sn alloy Stan et al., (2020). (e) A TEM image of silica particles in PMMA. (f) An SEM image of dual-phase steel Banerjee et al., (2013).
@@
Characterizing material microstructures has long been a major research focus. Recently, machine learning (ML), deep learning (DL), and AI have received growing interest in this area Stan et al., (2020); Maik Jablonka et al., (2020); Ajioka et al., (2020); Agrawal and Choudhary, (2019). These approaches transform traditional methods of characterization, enabling data-driven Processing-Structure-Properties-Performance (PSPP) linkages Agrawal and Choudhary, (2016). Traditionally, methods for pre-processing (for many downstream tasks) and analyzing multiphase micrographs were largely manual or semi-automated. For example, some methods rely on handcrafted descriptors (e.g., volume fraction, dispersion) Ashby and Cebon, (1993); Xu et al., (2014). Though intuitively appealing and sometimes effective, they are not generic to materials and fail to fully represent and distinguish complex morphologies Kondo et al., (2017). Statistical encoders (e.g., N-point correlation functions) can better represent microstructures but are often computationally expensive, depend on prior knowledge, and are limited in characterizing complex patterns in generic materials microstructures DeCost and Holm, (2015); Yabansu et al., (2017); Xu et al., (2015); Yu et al., (2017). The rich micrograph data sources that are now available warrant more automated methods that take advantage of modern machine learning algorithms. ML/DL can learn features directly from raw images. For instance, convolutional neural networks (CNNs) eliminate manual feature extraction and have achieved high accuracy in segmentation, phase identification, and particle size estimation DeCost et al., (2019); Jang et al., (2019); DeCost et al., (2017a). Even though many machine learning methods have been effectively applied to materials science problems, including micrograph segmentation, there remain challenges in expanding this interdisciplinary research: Much of the existing work centers on supervised learning that requires a large amount of labeled data (e.g., micrographs labeled pixel-by-pixel according to their microstructure class) Kondo et al., (2017); Lubbers et al., (2017); Azimi et al., (2018); Strohmann et al., (2019); Chen et al., (2019a) to train a segmentation model to reach a high accuracy. Moreover, works utilizing unsupervised or semi-supervised learning methods usually do not require labels during training Liu et al., (2024); Zhang et al., (2023); Huang et al., (2023) but still need manual identification and/or classification and are more expensive and less powerful than their supervised counterparts in production DeCost et al., (2017a); Impoco and Tuminello, (2015); Kitahara and Holm, (2018); Zhang et al., (2021). This study proposes a hybrid framework combining aspects of both approaches. By leveraging minimal labels and iterative refinement, we aim to produce accurate, scalable, and efficient microstructure characterization. In Section 2, we discuss the theoretical foundations and methods for each step of our framework. In Section 3, data sets, demonstrations, and experiment results for each step are presented, demonstrating that our approach can enhance characterization tasks with minimal human intervention.
@@ 2 Theoretical Foundation and Methodology
Figure 2: Flow chart of the iterative framework. Between steps, the knowledge passed from the previous to the next are shown: from Step $1$ to $2$, the HRs are extracted and entered into the database of existing MCs; from Step $2$ to $3$, the database of classified and labeled homogeneous microstructures is passed for data augmentation and training segmentation networks.
@@
Iteration over Steps $1$-$3$: As new micrographs are collected, they pass through the above three steps to classify/segment their phases and label any new, previously unseen, phases as new MCs (as in Step $2$) and to provide additional samples of existing MCs. The iterations are intended to improve the data-driven models (e.g., classification or segmentation neural networks) by either increasing the number of MCs they can classify (if one of the phases is labeled as a new MC) or by further training the models on the newly acquired additional samples of existing MCs.
@@ 2.1 Step 1: Micrograph nonstationarity analysis and unsupervised segmentation
Figure 3: Illustration of processing score vectors. (a) Neighborhood window (brown dashed square excluding the target pixel colored as red) for modeling the conditional distribution of $X_{i}$ (the red pixel) given the neighboring pixels in the window. (b) Spatial WMA window (blue square including the target pixel colored as red) for spatially smoothing score vectors. (c) A $2$D Gaussian weighting function is centered (at the red pixel) and truncated over the WMA window. The height of profile and grayscale value is proportional to the density value of the weighting function.
Any appropriate supervised learning model (e.g., a neural network) can be fit to all the observations $\{(X_{i},\bm{\mathbb{N}}({X}_{i}))\}_{i=1}^{m}$ (with $X_{i}$ and $\bm{\mathbb{N}}({X}_{i})$ being the response and the predictor variables, respectively) from a training micrograph(s), as a means of modeling the distribution $P({X}|\bm{\mathbb{N}}(X);\bm{\theta})$ with model parameters denoted by $\bm{\theta}$ (e.g., the weights for all neurons in a neural network). $\bm{\theta}$ implicitly represents the distribution of the pixels in the micrograph and therefore provides a relatively compact “fingerprint” that represents the stochastic nature of the microstructure. We adopt the approach developed in Zhang et al., (2021) for monitoring for nonstationarity in the nature of the microstructure using Fisher score vectors, which, for the $i^{\mathrm{th}}$ pixel, is defined as
@@
where $E_{\bm{\theta}}[\cdot]$ denotes the expectation operator (with respect to the random variables $X$ and $\bm{\mathbb{N}}(X)$), and the subscript $\bm{\theta}$ indicates that the distribution of pixel value $X$ and its neighborhood values $\bm{\mathbb{N}}(X)$ are based on the supervised learning model with true parameters $\bm{\theta}$. Conversely, if the local stochastic nature in the vicinity of pixel $X$ changes from what it was over the training data (i.e., the true parameters change from $\bm{\theta}$ to $\bm{\theta}^{\prime}\neq\bm{\theta}$ at $X$), then the mean of the score vector at $X$ for a model with parameters $\bm{\theta}$ generally differs from zero (i.e., $E_{\bm{\theta}^{\prime}}[\bm{s}(\bm{\theta};X,\bm{\mathbb{N}}(X))]\neq\bm{0}$). In practice, $\bm{\theta}$ for the training data is replaced by its maximum likelihood estimate $\hat{\bm{\theta}}$, and the expectation in Eq. (2) is replaced by the empirical mean $\hat{E}_{\bm{\theta}}[\bm{s}(\hat{\bm{\theta}};X,\bm{\mathbb{N}}(X))]$ over all pixels in the micrograph training data (for details refer to Zhang et al., (2021)). Here, the subscript $\bm{\theta}$ indicates whatever parameters are valid in the vicinity of pixel $X$, and $\hat{\bm{\theta}}$ indicates the globally estimated parameters over the training data.
From the preceding arguments, the problem of analyzing nonstationarity in the microstructure over a micrograph(s) is converted to fitting a single supervised learning model to the micrograph(s), applying the model to predict each pixel, and then analyzing whether the mean score vector locally differs substantially from the zero vector anywhere in the micrograph(s). If the stochastic nature of different MCs, as captured by $\bm{\theta}$, is distinct enough, the score vectors belonging to different HRs should center around different mean vectors and be separable in the score-vector space. Thus, a clustering algorithm can be applied to the (smoothed version of, via some spatial moving window) score vectors $\bm{s}(\hat{\bm{\theta}};y_{i},\bm{x}_{i})$ to diagnose the spatial nonstationarity of micrographs or, more specifically, segment different HRs based on their stochastic natures in an unsupervised manner Zhang et al., (2021). To mitigate the effects of noise in the score vectors on the clustering performance, before clustering, we calculate a spatial Gaussian weighted moving average (WMA) of score vectors, denoted as $\bm{z}_{i}$ for the $i^{\mathrm{th}}$ pixel, where the WMA window and the $2$D weight density functions are depicted in Fig. 3(b) and 3(c).
@@
$\mathcal{N}(\bm{\mu}_{k}|\bm{\mu}_{0},(\beta_{0}\bm{\Lambda}_{k})^{-1})$ is the multivariate Gaussian distribution with mean vector $\bm{\mu}_{0}$ and precision matrix $\beta_{0}\bm{\Lambda}_{k}$; and $\text{Wishart}(\bm{\Lambda}_{k}|\bm{W}_{0},\nu_{0})$ is the Wishart distribution with scale matrix $\bm{W}_{0}$ and degrees of freedom $\nu_{0}$. We follow standard rules Bishop, (2006); Blei et al., (2003); Teh et al., (2004) for choice of all hyperparameters $\bm{\alpha}_{0}$, $\bm{\mu}_{0}$, $\beta_{0}$, $\bm{W}_{0}$, and $\nu_{0}$. In particular, we use $\alpha_{0,k}=10^{-9}(k=1,2,\cdots,K)$, $\bm{\mu}_{0}=\frac{1}{m}\sum_{i=1}^{m}\bm{z}_{i}$, $\beta_{0}=1$, $\bm{W}_{0}^{-1}=\frac{1}{m-1}\sum_{i=1}^{m}(\bm{z}_{i}-\bar{\bm{z}})(\bm{z}_{i}-\bar{\bm{z}})^{T}$, and $\nu_{0}=dim(\bm{z}_{i})$. We then use the standard variational inference algorithm to compute posterior distributions for parameters $\bm{\pi}$, $\{\bm{\mu}_{k}\}_{k=1}^{K}$, and $\{\bm{\Lambda}_{k}\}_{k=1}^{K}$ and latent variables $\{\bm{c}_{i}\}_{i=1}^{m}$ and use their maximum a posteriori estimators as point estimates. The posterior estimate of the mixing proportion $\bm{\pi}$ can be used to determine the number of clusters in the micrograph, because the $k^{\mathrm{th}}$ element $\pi_{k}$ can be thought of as the portion of the effective number of $\bm{z}_{i}$’s assigned to the $k^{\mathrm{th}}$ cluster. This is a key feature of BGM that can help determine the number of clusters in a data-driven manner. See the later examples (e.g., Fig. 5(c) and 5(d)) for use of BGM posterior weights and AIC/BIC to determine the number of clusters. After estimating the number of clusters and clustering the score vectors based on the posterior distribution of $\{\bm{c}_{i}\}_{i=1}^{m}$, (each cluster represents a different MC) HRs can be obtained, possibly with minimal guidance from human experts to verify the unsupervised segmentation of HRs.
Note that over a region near the micrograph borders with width $l_{s}+l_{w}$ ($l_{s}$ and $l_{w}$ are the two window half-lengths in Fig. 3(a) and 3(b)), the pixels within this border region have no smoothed score vectors due to the lack of neighboring pixels and therefore cannot be segmented in Step $1$ (Fig. 6). Step $3$ (Sec. 2.3) avoids this limitation and can classify the MC in the region near borders by training a supervised segmentation model incorporating the MCs and verified labels from Steps $1$-$2$.

### v2

@@ Abstract
Microstructure of materials is often characterized through image analysis to understand processing-structure-properties linkages. We propose a largely automated framework that integrates unsupervised and supervised learning methods to classify micrographs according to microstructure phase/class and, for multiphase microstructures, segments them into different homogeneous regions. With the advance of manufacturing and imaging techniques, the ultra-high resolution of imaging that reveals the complexity of microstructures and the rapidly increasing quantity of images (i.e., micrographs) enables and necessitates a more powerful and automated framework to extract materials characteristics and knowledge. The framework we propose can be used to gradually build a database of microstructure classes relevant to a particular process or group of materials, which can help in analyzing and discovering/identifying new materials. The framework has three steps: (1) preliminary segmentation of multiphase micrographs so that different microstructure homogeneous regions can be identified in an unsupervised manner; (2) identification and classification of homogeneous regions of micrographs through an uncertainty-aware supervised classification network trained using the segmented micrographs from Step $1$ with their identified labels verified via the built-in uncertainty quantification and minimal human inspection; (3) subsequent supervised segmentation (more powerful than the segmentation in Step $1$) of multiphase microstructures through a segmentation network trained with micrographs and the results from Steps 11-22 using a form of data augmentation. This framework can iteratively characterize/segment new homogeneous or multiphase materials while expanding the database to enhance performance. The framework is demonstrated on various sets of materials and texture images.
@@ 1 Motivation and Introduction
Microstructure plays a vital role in determining physical properties of materials and understanding the relations among the processes, structures, and properties DeCost and Holm, (2015). Recently, analyzing images of multiphase materials, often called micrographs O’Mara et al., (2016); Puchala et al., (2016); Michel and Meredig, (2016), has attracted much attention across different materials fields (e.g., metals Lewandowski and Seifi, (2016), polymer composites Ligon et al., (2017), and ceramics Chen et al., (2019b); Moritz and Maleksaeedi, (2018), etc.) due to burgeoning manufacturing and imaging techniques Balla et al., (2012); Bandyopadhyay and Heer, (2018). For example, the ultra-high carbon steel (UHCS) data set DeCost et al., (2017b) includes $961$ SEM micrographs, of which two multiphase microstructures are shown in Figs. 1(a) and 1(b). Fig. 1(a) is a micrograph containing a single spheroidite phase, while Fig. 1(b) contains ferritic, proeutectoid, spheroidite, and Widmanstätten phases. Because the phase characteristics impact material properties (e.g., toughness and crack initiation etc.) Hecht et al., (2016); Carter et al., (2014), identifying and classifying the individual phases present in a micrograph is an important step towards understanding the physical properties of the material. Throughout this paper, by “multiphase” we mean a microstructure whose stochastic nature (different phases have different stochastic nature) varies spatially, sometimes referred to as microstructure nonstationarity. This includes certain steels that have different phases present simultaneously by design, as well as certain composites whose stochastic microstructure composition varies spatially due to variation in the processing conditions. By “single-phase”, we mean a homogeneous microstructure whose stochastic nature is spatially stationary. Throughout the paper, the term “microstructure class” (MC) refers to the label assigned to each homogeneous region (HR) of connected pixels within a micrograph with each HR comprised of a single phase.
@@
Figure 1: Examples of single phase (a) and multiphase materials (b)-(f)00footnotemark: 0. (a)-(b) Two SEM images of the UHCS data set DeCost et al., (2017b); Hecht et al., (2016). Notice that the two images are not in the same magnification, a commonly encountered issue of data quality in real data sets. (c) A X-ray CT (XCT) image of dendrite growth from a solidification experiment on a Al-Zn alloy Stan et al., (2020). (d) A serial sectioning ($S$) image of dendrite growth from a coarsening experiment on a Pb-Sn alloy Stan et al., (2020). (e) A TEM image of silica particles in PMMA. (f) An SEM image of dual-phase steel Banerjee et al., (2013).
@@
Characterizing material microstructures has long been a major research focus. Recently, machine learning (ML), deep learning (DL), and AI have received growing interest in this area Stan et al., (2020); Maik Jablonka et al., (2020); Ajioka et al., (2020); Agrawal and Choudhary, (2019). These approaches transform traditional methods of characterization, enabling data-driven Processing-Structure-Properties-Performance (PSPP) linkages Agrawal and Choudhary, (2016). Traditionally, methods for pre-processing (for many downstream tasks) and analyzing multiphase micrographs were largely manual or semi-automated and time consuming. For example, one week of synchrotron beamtime can produce $\sim 2.3$ millions micrographs, one of which is shown in Fig. 1(c), and manual segmentation of different metal phases of one micrograph takes a researcher $\sim 20$ minutes, which translates to $\sim 88$ years of human involvement to manually process one-week worth of data Stan et al., (2020). Another drawback of traditional methods is that they are not generic to materials and fail to fully represent and distinguish complex morphologies Kondo et al., (2017). For example, some methods rely on handcrafted descriptors (e.g., volume fraction, dispersion) Ashby and Cebon, (1993); Xu et al., (2014). Statistical encoders (e.g., N-point correlation functions (NPCF)) can better represent microstructures but are often computationally expensive, depend on prior knowledge of patch size, and are limited in characterizing complex and stochastic patterns in generic materials microstructures DeCost and Holm, (2015); Yabansu et al., (2017); Xu et al., (2015); Yu et al., (2017); Zhang et al., (2021). The rich micrograph data sources that are now available warrant more automated methods that take advantage of modern machine learning algorithms. ML/DL can learn features directly from raw images. For instance, convolutional neural networks (CNNs) eliminate manual feature extraction and have achieved high accuracy in segmentation, phase identification, and particle size estimation DeCost et al., (2019); Jang et al., (2019); DeCost et al., (2017a). Even though many machine learning methods have been effectively applied to materials science problems, including micrograph segmentation, there remain challenges in expanding this interdisciplinary research: Much of the existing work centers on supervised learning that requires a large amount of labeled data (e.g., micrographs labeled pixel-by-pixel according to their microstructure class) Kondo et al., (2017); Lubbers et al., (2017); Azimi et al., (2018); Strohmann et al., (2019); Chen et al., (2019a) to train a segmentation model to reach a high accuracy. Moreover, works utilizing unsupervised or semi-supervised learning methods usually do not require labels during training Liu et al., (2024); Zhang et al., (2023a); Huang et al., (2023) but still need manual identification and/or classification and are more expensive and less powerful than their supervised counterparts in production Zhang et al., (2021); DeCost et al., (2017a); Impoco and Tuminello, (2015); Kitahara and Holm, (2018). This study proposes a hybrid framework combining aspects of both approaches. By leveraging minimal labels and iterative refinement, we aim to produce accurate, scalable, and efficient microstructure characterization.
A main contribution of this paper is the overall framework that integrates various unsupervised learning and supervised learning concepts to initially segment a micrograph into HRs (using unsupervised learning), classify the identified HRs, and then refine the segmentation using a supervised segmentation approach. Applied iteratively to a series of micrographs, the framework can be used to build a database of stochastic microstructures that is incorporated into the supervised segmentation and MC classification steps. The unsupervised learning Step $1$ is based on the approach of Zhang et al., (2021) but uses Bayesian Gaussian Mixture clustering to more systematically determine the number of HRs. The classification Step $2$ is novel in the sense that it is first informed by the unsupervised segmentation from Step $1$ so that each identified HR from Step $1$ is classified as a single entity (as opposed to classifying each pixel or small region of pixels in the original micrograph), which is more efficient since it allows information over the entire HR to be used for classification. We also incorporate evidential deep learning into Step $2$, which allows the framework to assess the likelihood that a newly collected HR belongs to a new MC that is not currently in the database. The supervised segmentation in Step $3$ incorporates a novel form of transfer learning using a database of stochastic texture images (of various fabrics, grass, seeds, ceiling textures, etc) to pretrain the model. Although these pretraining images are not of material microstructures, we demonstrate that it substantially improves the material microstructure segmentation and classification and mitigates the problem of collecting labelled segmentation data of materials microstructures.
The format of the remainder of the paper is as follows. In Section 2, we discuss the theoretical foundations and methods for each step of our framework. In Section 3, data sets, demonstrations, and experiment results for each step are presented, demonstrating that our approach can enhance characterization tasks with minimal human intervention.
@@ 2 Theoretical Foundation and Methodology
Figure 2: Flow chart of the iterative framework. Between steps, the knowledge passed from the previous to the next are shown: from Step 11 to $2$, the HRs are extracted and entered into the database of existing MCs; from Step 22 to $3$, the database of classified and labeled homogeneous microstructures is passed for data augmentation and training segmentation networks.
@@
Iteration over Steps 11-33: As new micrographs are collected, they pass through the above three steps to classify/segment their phases and label any new, previously unseen, phases as new MCs (as in Step $2$) and to provide additional samples of existing MCs. The iterations are intended to improve the data-driven models (e.g., classification or segmentation neural networks) by either increasing the number of MCs they can classify (if one of the phases is labeled as a new MC) or by further training the models on the newly acquired additional samples of existing MCs.
@@ 2.1 Step 1: Micrograph nonstationarity analysis and unsupervised segmentation
Figure 3: Illustration of processing score vectors. (a) Neighborhood window (brown dashed square excluding the target pixel colored as red) for modeling the conditional distribution of $X_{i}$ (the red pixel) given the neighboring pixels in the window. (b) Spatial WMA window (blue square including the target pixel colored as red) for spatially smoothing score vectors. (c) A 22D Gaussian weighting function is centered (at the red pixel) and truncated over the WMA window. The height of profile and grayscale value is proportional to the density value of the weighting function.
Any appropriate supervised learning model (e.g., a neural network) can be fit to all the observations $\{(X_{i},\bm{\mathbb{N}}({X}_{i}))\}_{i=1}^{m}$ (with $X_{i}$ and $\bm{\mathbb{N}}({X}_{i})$ being the response and the predictor variables, respectively) from a training micrograph(s), as a means of modeling the distribution $P({X}|\bm{\mathbb{N}}(X);\bm{\theta})$ with model parameters denoted by $\bm{\theta}$ (e.g., the weights for all neurons in a neural network). $\bm{\theta}$ implicitly represents the distribution of the pixels in the micrograph and therefore provides a relatively compact “fingerprint” that represents the stochastic nature of the microstructure. We adopt the approach developed in Zhang et al., (2021) for monitoring for nonstationarity in the nature of the microstructure using Fisher score vectors (which was based on the more general concept drift detection approach of Zhang et al., (2023b)), which, for the $i^{\mathrm{th}}$ pixel, is defined as
@@
where $E_{\bm{\theta}}[\cdot]$ denotes the expectation operator (with respect to the random variables $X$ and $\bm{\mathbb{N}}(X)$), and the subscript $\bm{\theta}$ indicates that the distribution of pixel value $X$ and its neighborhood values $\bm{\mathbb{N}}(X)$ are based on the supervised learning model with true parameters $\bm{\theta}$. Conversely, if the local stochastic nature in the vicinity of pixel $X$ changes from what it was over the training data (i.e., the true parameters change from $\bm{\theta}$ to $\bm{\theta}^{\prime}\neq\bm{\theta}$ at $X$), then the mean of the score vector at $X$ for a model with parameters $\bm{\theta}$ generally differs from zero (i.e., $E_{\bm{\theta}^{\prime}}[\bm{s}(\bm{\theta};X,\bm{\mathbb{N}}(X))]\neq\bm{0}$). In practice, $\bm{\theta}$ for the training data is replaced by its maximum likelihood estimate $\hat{\bm{\theta}}$, and the expectation in Eq. (2) is replaced by the empirical mean $\hat{E}_{\bm{\theta}}[\bm{s}(\hat{\bm{\theta}};X,\bm{\mathbb{N}}(X))]$ over all pixels in the micrograph training data (for details refer to Zhang et al., (2021)). Here, the subscript $\bm{\theta}$ indicates whatever parameters are valid in the vicinity of pixel $X$, and $\hat{\bm{\theta}}$ indicates the globally estimated parameters over the training data. The estimate $\hat{\bm{\theta}}$ fit to the entire micrograph can be thought of as an aggregation of the $\bm{\theta}$ for each MC in the micrograph, and the average of the score vectors $\bm{s}(\hat{\bm{\theta}};y_{i},\bm{x}_{i})$ over the entire micrograph is the zero vector.
From the preceding and the following arguments, which summarize the theory in Zhang et al., (2021), the problem of analyzing nonstationarity in the microstructure over a micrograph(s) is converted to fitting a single supervised learning model to the micrograph(s), applying the model to predict each pixel, and then analyzing whether the mean score vector locally differs substantially from the zero vector anywhere in the micrograph(s). If the stochastic nature of different MCs, as captured by $\bm{\theta}$, is distinct enough, the score vectors $\bm{s}(\hat{\bm{\theta}};y_{i},\bm{x}_{i})$ belonging to different HRs should center around different mean vectors and be separable in the score-vector space. Thus, a clustering algorithm can be applied to the (smoothed version of, via some spatial moving window) score vectors to diagnose the spatial nonstationarity of micrographs or, more specifically, segment different HRs based on their stochastic natures in an unsupervised manner Zhang et al., (2021). To mitigate the effects of noise in the score vectors on the clustering performance, before clustering, we calculate a spatial Gaussian weighted moving average (WMA) of score vectors, denoted as $\bm{z}_{i}$ for the $i^{\mathrm{th}}$ pixel, where the WMA window and the 22D weight density functions are depicted in Figs. 3(b) and 3(c).
@@
$\mathcal{N}(\bm{\mu}_{k}|\bm{\mu}_{0},(\beta_{0}\bm{\Lambda}_{k})^{-1})$ is the multivariate Gaussian distribution with mean vector $\bm{\mu}_{0}$ and precision matrix $\beta_{0}\bm{\Lambda}_{k}$; and $\text{Wishart}(\bm{\Lambda}_{k}|\bm{W}_{0},\nu_{0})$ is the Wishart distribution with scale matrix $\bm{W}_{0}$ and degrees of freedom $\nu_{0}$. We follow standard rules Bishop, (2006); Blei et al., (2003); Teh et al., (2004) for choice of all hyperparameters $\bm{\alpha}_{0}$, $\bm{\mu}_{0}$, $\beta_{0}$, $\bm{W}_{0}$, and $\nu_{0}$. In particular, we use $\alpha_{0,k}=10^{-9}(k=1,2,\cdots,K)$, $\bm{\mu}_{0}=\frac{1}{m}\sum_{i=1}^{m}\bm{z}_{i}$, $\beta_{0}=1$, $\bm{W}_{0}^{-1}=\frac{1}{m-1}\sum_{i=1}^{m}(\bm{z}_{i}-\bar{\bm{z}})(\bm{z}_{i}-\bar{\bm{z}})^{T}$, and $\nu_{0}=dim(\bm{z}_{i})$. We then use the standard variational inference algorithm to compute posterior distributions for parameters $\bm{\pi}$, $\{\bm{\mu}_{k}\}_{k=1}^{K}$, and $\{\bm{\Lambda}_{k}\}_{k=1}^{K}$ and latent variables $\{\bm{c}_{i}\}_{i=1}^{m}$ and use their maximum a posteriori estimators as point estimates. The posterior estimate of the mixing proportion $\bm{\pi}$ can be used to determine the number of clusters in the micrograph, because the $k^{\mathrm{th}}$ element $\pi_{k}$ can be thought of as the portion of the effective number of $\bm{z}_{i}$’s assigned to the $k^{\mathrm{th}}$ cluster. This is a key feature of BGM that can help determine the number of clusters in a data-driven manner. See the later examples (e.g., Figs. 5(c) and 5(d)) for use of BGM posterior weights and AIC/BIC to determine the number of clusters. After estimating the number of clusters and clustering the score vectors based on the posterior distribution of $\{\bm{c}_{i}\}_{i=1}^{m}$, (each cluster represents a different MC) HRs can be obtained, possibly with minimal guidance from human experts to verify the unsupervised segmentation of HRs.
Note that over a region near the micrograph borders with width $l_{s}+l_{w}$ ($l_{s}$ and $l_{w}$ are the two window half-lengths in Figs. 3(a) and 3(b)), the pixels within this border region have no smoothed score vectors due to the lack of neighboring pixels and therefore cannot be segmented in Step $1$ (Fig. 6). Step $3$ (Sec. 2.3) avoids this limitation and can classify the MC in the region near borders by training a supervised segmentation model incorporating the MCs and verified labels from Steps 11-22.